OpenStudy (anonymous):

The sum of a number and it's reciprocal is 13 over 6. Find the number.

OpenStudy (anonymous):

@anonymous_user

OpenStudy (anonymous):

so The sum of a number and it's reciprocal is \[\frac{ 13 }{ 6 }\]. Find the number.

OpenStudy (anonymous_user):

2/3

OpenStudy (anonymous):

or in the Language of Algebra it would be \[n +\frac{ 1 }{ n }= \frac{ 13 }{ 6 }\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous_user):

De nada